40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 185 of 341  Not logged in ELOG logo
ID Date Author Type Category Subject
  7870   Fri Dec 21 19:49:39 2012 RijuUpdate Photodiode transimpedance

I have repeated the transimpedance measurement of PDA10CF. Also made the dark current noise measurement by connecting the PDA10CF output to the A channel of network analyzer.  The results are as follows. I I started to take the reading for shot noise intercept current using a light bulb in front of the PD, changing the current through the bulb, but at higher current the bulb filament got broken, so the experiment is incomplete.

  7869   Fri Dec 21 16:50:30 2012 RanaUpdateSUSTT in vac DB25 pin swapping

[Koji, Rana, Nic, Steve]

We went to the 25-pin D cable which connects to the TT1 quadropus and succeeded eventually in swapping pins 12/24 into the 13/25 positions.

  1. The D-sub connector is a custom made LIGO part and so it doesn't at all work to use the standard pin extractor tools to move the pins out; we should have investigated this before spending all this time poking at and possibly damaging the existing connector.
  2. To move the pins, we have to partially dis-assemble the connector and fish the pins/wires through the appropriate holes. Unfortunately, the design is such that we nearly lose all of the pins when trying to do this. Pictures describe the story better than words.
  3. After the swap we tried to test the TT, but again wasted some time because the vac feedthrough was incorrectly labeled. The 25-pin feedthrough labeled as "PZT1" does not, in fact, connect to the TT. Instead, its the one slightly above it that is labeled "Pico". I have moved the PZT1 sticker up to match the actual connector. In order to discover this, we beeped through several stages of the coil driver, cable system. WE need to order some in-line D-sub breakouts for 25pin, 37pin, and 9pin which are similar to the ones we have now for 15pin. These are better than the green terminal block breakouts.
  4. After this, we were able to see the TT move, but elected to leave the final piece of the work (determining which microD goes with which coil) to when Jamie gets back.
  5. The TT screen is not good: it needs to be just like the usual sus screen so that we can put in offsets, excitations, etc. Perhaps also the ASC-TT screen can link to the TT:SUS screens. We can just copy the eLIGO TT screens to get going.
  7868   Fri Dec 21 09:55:23 2012 SteveUpdateGeneralin-vac adapter cables for TTs

 

 Accu-Glass is closed till Jan 2, 2013

Atm2 would be the ideal solution Gender Adaptor Special with male - female sides and vented D-sub. How many are we getting 2 or 6 ?

  7866   Thu Dec 20 19:46:20 2012 ranaConfigurationEnvironmentControl Room Projector

 Needs

StevePurple.JPG

  7865   Thu Dec 20 18:33:52 2012 JamieUpdateGeneralin-vac adapter cables for TTs

We need short cables that mirror the pins:

invac-adapter-cable.pdf

The male side will plug into the 25-pin female on the stack-top bracket.  The tip-tilt quadrapus cable will plug into the female side. This will match up pin 1 on the tip-tilt cable, which is connected to it's shield, to pin 13 on the bracket, which is the shield of the cable that runs to the stack.

They need to be vacuum compatible.  Shorter length is preferred, and there is no minimum length (something like an all-in-one gender changer would be ideal, but probably expensive to have made).


 

  7864   Thu Dec 20 17:13:56 2012 JamieUpdateGeneralhow to deal with problem with in-vac wiring for TTs

So this is obviously a general problem for all the TTs.  Our in-vacuum wiring is unfortunately mirrored relative to that of aLIGO, or at least:

  • aLIGO: in-vacuum pin 1 tied to shield (T1200131)
  • 40m: in vacuum pin 13 tied to shield

And again, the problem is that pin 13 on the TT quadrapus cables is the one of the coil pins for one of the OSEMs.

I think the right solution is to make mirroring adapter cables for the TTs.  Modifying the pins on the stack-top brackets for just the TTs would leave us with a bunch of brackets that are different than all the rest, which I think is a bad idea.  Therefore we leave all the feed-thru-->bracket wiring the same, and make adapters.  I'll describe the adapters in a follow-up post.

The silver lining to this whole thing, if there is one, is that I wired the polarity on the out-of-vac adapter cable at the coil driver in such a way that the drive/send signal went to the grounded in-vac pin.  If I had by chance wired the polarity oppositely everything would probably have worked, except that the return for one of the coils would have gone through the cable shield and the chamber, rather than the return pin to the coil driver.  I'll let you image the problems that would have caused.

Making new adapters will take a little while, but I think we can proceed with the installation and alignment with a temporary setup in the mean time by taking advantage of the polarity I mention above.  We can temporarily swap the polarity so that we can drive current to the coil using pin 13.  This will allow us to complete the installation and do all the alignment.  Once the in-vac adapter cable arrives, we just put it in, fix the out-of-vac polarity, check that everything works as expected, and button up.

We'll pick up all this when we're back on Jan 7.  Steve will put in an order for the in-vac adapter cables ASAP.

I take full responsibility for this fuck up.  We've been unable to find any in-vac wiring diagrams, but I should have checked all of the wiring during the last vent so that we could have prepared for this ahead of time.  Sorry.

  7863   Thu Dec 20 12:14:19 2012 JamieUpdateGeneralproblem with in-vac wiring for TTs

Nic and I discovered a problem with the in-vac wiring from the feed-thru to the top of the table.  Pin 13 at the top of the stack, which is one of the coil pins on the tip-tilt quadrapus cables, is *the* shield braid on the cable that goes to the feed-thru.  This effectively shorts one of our coil signals.

There are three solutions as we see it:

* swap pin 13 for something else at the top of the stack, and then swap it back somewhere else outside of the vacuum.

* swap *all* the pins at the top of the table to be the mirror.  We would then need to mirror our cables on the outside, but that's less of an issue.

* make a mirror adapter that sits at the top.  This would obviously need to be cleaned/baked.

None of these solutions is particularly good or fast.

  7862   Thu Dec 20 10:35:27 2012 SteveUpdateIOOTT1 connections

Quote:

[Jamie,Manasa]

We've been trying to figure out the connector for the TTs. Since, we found the cables were plugged in wrong in TT1; when triggered in pitch, the mirror moves in yaw and viceversa.

Referring to the cabling diagram, D1000234-v10, we infer that connectors go as J2 - LR, J3 - UR, J4 - LL, J5 - UL and the connections are made looking at the mirror from behind.

 The view looking at the optic from the back:  

UL    UR

LL    LR

 

 

  7861   Thu Dec 20 10:11:12 2012 ManasaUpdateIOOTT1 connections

[Jamie,Manasa]

We've been trying to figure out the connector for the TTs. Since, we found the cables were plugged in wrong in TT1; when triggered in pitch, the mirror moves in yaw and viceversa.

Referring to the cabling diagram, D1000234-v10, we infer that connectors go as J2 - LR, J3 - UR, J4 - LL, J5 - UL and the connections are made looking at the mirror from behind.

  7860   Wed Dec 19 21:35:33 2012 KojiSummaryGeneralFirst Contact Training with Margot

[Koji, Steve]

First Contact Training with Margot

  7859   Wed Dec 19 20:18:51 2012 ranaUpdateComputersWe are Changing the Passwerdz next week----

Be Prepared

http://xkcd.com/936/

  7858   Wed Dec 19 19:28:12 2012 ManasaUpdateIOOPZT1 removed, TT1 in place

Quote:

Tip Tilt, serial number ### (Manasa will get the serial number and put it in the elog) was taken out of the cleanroom, for use as TT1.

 

Depending on how far things get tonight, Jamie and Manasa may ask Steve to help them remove the BS door, so they can get started on replacing PZT2 with TT2.

Tip-Tilt TT1

 

I have fixed TT1 close to what it's position looks like in the CAD drawing. Only 2/3 of TT1 rests on the table...so we need to be extra careful when we will move it for alignment.

Serial Number:  SN 027

dcc number: D1001450-V2

 We are still in the process of removing the quadrupus from the TT, figuring out which connector goes where, putting it on correctly, and re-testing.

We closed the IMC chamber with light doors calling it a day!

 

  7857   Wed Dec 19 18:40:00 2012 JenneUpdateIOOPZT1 removed, TT1 in place

[Manasa, Jamie, Jenne]

PZT1 has been removed, and is wrapped in foil and stored in a (labeled) plastic box.

We beeped the cable between the cable holder bracket on the in-vac table, and the outside of the feedthrough.  Things are mirrored, so pins 1,14 (one edge on the feedthrough) go to pins 13,25 on the in-vac cable bracket.

Tip Tilt, serial number ### (Manasa will get the serial number and put it in the elog) was taken out of the cleanroom, for use as TT1.

We checked the epics controls from the TT screen that Jamie made a while back (accessible from the ASC tab on the sitemap) to the output of the AI board.  Things were very weird, but Jamie fixed them up in the model, then rebuilt and restarted the ASS model so that now the epics channel corresponding to, say, UL actually actuates on the UL output of the boards.

We tested the cables from the rack to the feedthrough, and discovered that they are also mirrored, to undo the mirroring between the feedthrough and the in-vac bracket.

Jamie made an adapter cable to take the pinout of the coil driver boards correctly to the pinout of the quadrupus cable, through this double-mirroring (i.e. no net mirror effect).

We set up a laser pointer on a tripod outside the door of the MC chamber (where the access connector usually is), and pointed it at the back of the TT.  Den or whomever put the cable on the TT didn't follow the diagram (or something got messed up somewhere), because when we actuate in pitch (+ on the uppers, - on the lowers), we see the TT move in yaw, and vice versa. 

We are in the process of removing the quadrupus from the TT, figuring out which connector goes where, putting it on correctly, and re-testing.

Depending on how far things get tonight, Jamie and Manasa may ask Steve to help them remove the BS door, so they can get started on replacing PZT2 with TT2.

  7856   Wed Dec 19 11:53:23 2012 SteveUpdateVACAC removed

 Koji, Manasa, Jenne, Jamie, Bob and Steve

Access connector removed this morning and work has began in the IOO chamber.  BE WARE OF ANTS !

 

  7855   Wed Dec 19 11:14:25 2012 SteveUpdateSUSBeCu wire in stock

Quote:

Just in case we want to retrofit the Tip/Tils with Beryllium Copper wire, here are links to a few sources which have a supply of the right composition and temper:

http://www.lfa-wire.com/Tempered-Alloy-25_C17200.htm

http://www.alloywire.com/beryllium_copper_CB_101.html

http://www.ngk.co.jp/english/products/electronics/berylliumcopper/wire/index.html

http://www.goodfellow.com/E/Copper-Beryllium-Wire.html

 

I don't think its worth it to do something to modify them unless we get a real reduction in the hysteresis - need a benchtop test setup ASAP.

 Be Copper in the lab is from Ca Fine Wire :  alloy 10 CDA 17 in sizes .008"  &  0.002"  There are other sus wire choices in the Drever lab

  7854   Tue Dec 18 16:44:00 2012 rijuUpdate Photodiode transimpedance

Today I measured the dark current of the PDA10CF. The output of the PD was connected to the A channel of the network analyzer, when there was no light falling on it. The response is collected using GPIB.

I will upload the result shortly.

  7853   Tue Dec 18 16:37:40 2012 SteveUpdateVAClast RGA scan before vent

 

 

  7852   Tue Dec 18 16:37:17 2012 JamieUpdateAlignmentPost vent, pre door removal alignment

[Jenne, Manasa, Jamie]

Now that we're up to air we relocked the mode cleaner, tweaked up the alignment, and looked at the spot positions:

mcspot_post_vent.pdf

The measurements from yesterday were made before the input power was lowered.  It appears that things have not moved by that much, which is pretty good.

We turned on the PZT1 voltages and set them back to their nominal values as recorded before shut-down yesterday.  Jenne had centered IPPOS before shutdown (IPANG was unfortunately not coming out of the vacuum).  Now we're at the following value: (-0.63, 0.66).  We need to calibrate this to get a sense of how much motion this actually is, but this is not insignificant.

 

  7851   Tue Dec 18 15:51:33 2012 JenneUpdateIOOOld G&H TT mirrors' phase maps measured

I took the 2 G&H mirrors that we de-installed from PR3 and SR3 over to GariLynn to measure their phase maps. Data is in the same place as before, http://www.ligo.caltech.edu/~coreopt/40MCOC/Oct24-2012/ .  Optic "A" is SN 0864, and optic "B" is SN 0884, however I'm not sure which one came from which tip tilt.  It's hard to tell from what photos we have on picasa.

Both are astigmatic, although not lined up with the axes defined by where the arrow marks the HR side.  Both have RoCs of -600 or -700m. RMS of ~10nm.

  7850   Tue Dec 18 15:32:29 2012 JenneUpdateSUSTT angle of incidence

Quote:

At least, we don't want to use Al-coated mirrors. We should use multilayer dielectric mirrors.

 I popped into the cleanroom earlier today, and all 4 active TTs have dielectric coatings.  I'm not sure why the mirror in this photo looks funny, but the actual mirrors installed are correct, at least in type of coating. 

I'm not sure if Den wrote down what mirrors are actually in there, and I didn't look carefully - I don't know if they are G&H, CVI, other mystery company?

  7849   Tue Dec 18 10:43:20 2012 SteveUpdateVACvent has completed

 The 40m vacuum envelope is at atm    It took 4.5 hours at ~3 Torr/min speed. Atm1

 

What I did:

Turned PZT voltages off: 84.1V at OMC-rack location

                                           -  99V at 1Y3 location as PZT_pitch,

                                             56.6V KEPCO at 1Y3 location as PZT_yaw                                           

Took screen shuts of SUS_SUMMARY, yesterday,  Atm3

Turned off OPLEV _SERVOs, today , Atm2

Closed PSL_SHUTTER and green shutters at the ends

 

Set vacuum valves for vent position and opened VM2 for the RGA

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

  7848   Tue Dec 18 07:03:40 2012 ManasaUpdateGeneralVent plan: WE VENT TOMORROW

Quote:

 

12/19: First thing in the morning we take off the access connector ONLY.  The access connector is all we'll need to replace PZT1.    Put light door on the OMC chamber side immediately, since we won't need to get in there at all.  We won't need the light access connector.

 The access connector is all we'll need to replace PZT1.

Really? Can you be sure of the input pointing this way?

 

 

 

 

 

 

  7847   Tue Dec 18 00:45:19 2012 KojiSummaryGeneralThe projector lamp ended its life?

...Nah. The projector is pretty new (t<1yr) and this is the first time to have the lamp busted after the installation last year in Jan.

We just should purchase two bulbs.

  7846   Mon Dec 17 22:49:33 2012 ManasaSummaryGeneralThe projector lamp ended its life?

Quote:

i just heard a rather large exploding sound in the control room.
I tried to locate the source and found the projector is not illuminating the wall anymore.
There is a slight smell of burning, but nothing is smoking.

Probably the lamp ended its life.

Rana and I just talked about the projector life time an hour ago! It must have been hearing!

 LOL

We should try purchase a projector with LED this time...longer lifetime! I guess the price of replacing the lamp in the one we have will be more or less same as a new one!

  7845   Mon Dec 17 22:42:27 2012 KojiSummaryGeneralThe projector lamp ended its life?

i just heard a rather large exploding sound in the control room.
I tried to locate the source and found the projector is not illuminating the wall anymore.
There is a slight smell of burning, but nothing is smoking.

Probably the lamp ended its life.

Rana and I just talked about the projector life time an hour ago! It must have been hearing!

  7844   Mon Dec 17 21:41:30 2012 ranaSummarySUSBeCu wire

Just in case we want to retrofit the Tip/Tils with Beryllium Copper wire, here are links to a few sources which have a supply of the right composition and temper:

http://www.lfa-wire.com/Tempered-Alloy-25_C17200.htm

http://www.alloywire.com/beryllium_copper_CB_101.html

http://www.ngk.co.jp/english/products/electronics/berylliumcopper/wire/index.html

http://www.goodfellow.com/E/Copper-Beryllium-Wire.html

 

I don't think its worth it to do something to modify them unless we get a real reduction in the hysteresis - need a benchtop test setup ASAP.

  7843   Mon Dec 17 21:33:45 2012 ManasaUpdateGeneralVent plan: WE VENT TOMORROW

Quote:

Manasa is lowering the input power and preparing the mode cleaner.

I aligned the PSL angle and position QPDs...then attenuated the input laser power using HWP-PBS-HWP combo from 1.28W to 100mW following instructions from elog 6892 and elog 7299.

To enable MC locking, I replaced the 10% BS before the MC_REFL PD with Y1 mirror. I tweaked the steering mirrors at the PSL table a teeny tiny bit to enable MC locking. MC is now locked at low power with 1.0 transmission.

  7842   Mon Dec 17 20:22:07 2012 JamieUpdateGeneralVent plan: WE VENT TOMORROW

We will vent tomorrow (12/18)

After lengthy discussion, I have determined that we should vent now, with the primary goal of this vent to replace the input steering PZTs with the new active tip-tilts.  Since we still don't understand exactly what's going on with the PRC, it's unclear what we would do to attack the problem.  We need to do more modeling and measurements first.  We should limit the goal of this vent to replacing the PZTs, and then close up and do more measurements with better modeling and improved input point in hand.

There is limited time this week before everyone leaves for two-week holidays on Thursday or Friday.  The reason to not vent would be that we don't want to leave the IFO at air during the two week holiday.  People seem to think that this is not a problem, so we don't gain anything by waiting.  Therefore we vent now and do what we can before people take off.

The goal for this week is to replace PZT1 only:

Now: Jenne and Manasa are doing vent prep.  Manasa is lowering the input power and preparing the mode cleaner.  Jenne centered IPPOS and IPANG.  This will allow us to check how the input alignment changed during the vent.

12/18: Steve is going to start the vent as soon as he gets back from the dentist, at around 10am.  He will regulate the vent such that we are ready to lock the mode cleaner by 4pm.  At that time we will lock the MC and recheck the input alignment with IPPOS/ANG, with the idea being to see if anything moves during the vent.

12/19: First thing in the morning we take off the access connector ONLY.  The access connector is all we'll need to replace PZT1.    Put light door on the OMC chamber side immediately, since we won't need to get in there at all.  We won't need the light access connector.

For the rest of Wednesday we'll remove PZT1 and install the new active TT.  We'll be using the current out-of-vac cable for PZT1 for the new TT.  We should only have to modify the rack end of the cable to accommodate the coil driver.  This should be a small modification.  Given that we have no wiring diagrams we'll have to pin it out in situ.

12/20: Hopefully finish up TT installation.

Jenne leaves 12/20, Manasa and Jamie leave 12/21.  We will either leave light doors on access connector holes, or possibly Rana, Koji, and Steve will replace access connector on Friday so that we can pump down to 1 Torr or something so that we leave it there over the holiday.

After we return from vacation:

PZT2/TT2 installation.  This will be less straight forward since the new TT has a bigger foot print than PZT2 and will block the PRM optical levers.  We'll need one additional steering mirror to redirect the oplev around the TT.  See elog 7815.

Once the new TTs are installed, we'll reevaluate where we're at.  If PRC modeling has progressed and we have an idea of something to work on with the PRC, we can.  Otherwise, we'll just button up, pump down, and get on with some better PRC measurements.

  7841   Mon Dec 17 19:47:15 2012 ranaUpdatePEMSeismic StripTool config updated

I have updated the Seismic Striptool display which is plotted on the wall in the control room. Please take a look and make comments. We should finalize it and not change it anymore.

By having an unchanging display, we can get used to small changes in the seismic environment which disrupt our locking.

  1. y-scale is now linear; the log-scale was suppressing the factor of 2-3 variations which are important to us.
  2. Just as the rainbow does, the colors now go from red to purple to represent the noise from 0.1 - 30 Hz: the red traces are 0.1-0.3 Hz, the green/yellow traces are 0.3-3 Hz, and the blue/purple traces are 3-30 Hz.
  3. This is just showing GUR1. Let's try to keep this seismometer working so that we can have some long term record of the seismicity here. This means don't click off the buttons, disconnect the sensor, reboot the machine, etc. When you do do these things, elog them.
  7840   Mon Dec 17 18:39:02 2012 JenneUpdateIOOIPPOS centered, no IPANG beam

I centered IPPOS.  I do not find any IPANG beam on the ETMY table, even waving the IR card in the black beam tube.

I took photos of the PZT2 HV supplies with the new camera, but I can't figure out how to get the pictures off the camera.  I guess the camera is smarter than I am.

  7839   Mon Dec 17 14:45:01 2012 JenneUpdateAlignmentVideos with PRMI locked

[Jamie, Jenne]

Koji and Jamie locked the PRMI, and then Jamie and I took some videos. 

Video 1:   https://www.youtube.com/watch?v=jszTeyETyxU shows the face of PR2.

Video 2:   https://www.youtube.com/watch?v=Tfi4I4Q3Mqw shows the back of PR3, the face of PR2, as well as REFL and AS.

Video 3:   https://www.youtube.com/watch?v=bLHNWHAWZBA is the camera looking at the face of PRM and (through a viewing mirror) BS.

 

If you watch video 1, you'll see how large the beam gets on the face of PR2.  The main spot, where the straight-through, no-cavity beam is, is a little high of center.  The rest of the inflated beam swirls around that point.

Video 2 shows the same behavior, but you also see that we're much too high on PR3, and too close to the right (as seen on the video) side.

Video 3 is very disconcerting to me.  The main, stationary beam spot seems nicely centered, but the resonant beam, since it inflates and gets big, is very close to the right side of the PRM (as seen on the video). 

It wouldn't surprise me if, were we able to quantify the beam clipping loss on PR3 and PRM, the clipping were the reason we have a crappy PRC gain.  This doesn't explain why we have such a weird inflated beam though.

  7838   Mon Dec 17 14:13:55 2012 JenneUpdateSUSPRM oplev gains restored

Quote:

PRM oplev gains set to zero from PIT 0.15 and YAW -0.3 and damping restored

 Put them back to normal.

  7837   Mon Dec 17 11:20:58 2012 JenneUpdateSUSBeam dumps on vertex oplevs removed

I'm not sure when this was done, but there were beam dumps in front of the lasers for BS/PRM oplevs as well as ITMY/SRM oplevs.  MICH wasn't holding lock very nicely, so I poked around, and the Sum values for all of these optics' oplevs seemed too low, so I went to look, and found dumps.  I have removed these, and now BS and ITMY oplevs are back to normal.  (PRM and SRM are still misaligned right now, so I'll check those later, but they should be fine).

BS's oplev has been enabled while non-existant, at least for the whole weekend, since I found it enabled.  ITMY I found misaligned, so it's oplev servos were off.

In other news, we should get back in the habit of restoring all optics before we leave for the night / whenever locking activities are finished. 

  7836   Fri Dec 14 17:12:19 2012 Evan HallUpdatePSLPMC yaw tune-up (from Wednesday night)

Wednesday night, there was ~0.4 V on the PMC transmission PD. I adjusted the steering mirrors into the PMC and got the transmission up to 0.81 V.

  7835   Fri Dec 14 16:35:38 2012 AyakaUpdateLSCHigh frequency sensitivity improved

Since I found that the the AS sensitivity seems to be limited by circuit noise, I inserted a RF amplifier just after the AS RF output.
Now, the sensitivity is improved and limited by the dark noise of the PD.

ASspe_noise_amp.pdf

(Note: I did not apply the open loop TF on this xml file.)
REF3: dark noise + circuit noise + WT filter noise + ADC noise
REF4: circuit noise + WT filter noise + ADC

With this situation, I injected the acoustic noise:

ASspe_acoustic_amp.pdf

REF5, 6, 7: with acoustic excitation
no reference: without acoustic excitation

We could see the coherence only at the same frequencies, around 200 Hz as we saw before (elog).

  7834   Fri Dec 14 14:40:31 2012 RijuUpdate Photodiode transimpedance

Photodiode PDA10CF was under test. The RF out signal of AG4395A had been divided by splitter with one output of the splitter going to R channel of the network analyzer and the other to the laser. The splitted laser beams - splitted with beam splitter - fall on two photodiodes - one reference and the other on PDA10CF. The outputs of these two photodiodes go to channel B and A respectively of the network analyzer. The measured transimpedance data had been collected using the GPIB connection.

The result is as follows:

  7833   Fri Dec 14 10:09:30 2012 AyakaUpdateLSCHigh frequency noise in AS signal

Quote:

This is NOT calibrated. Its sort of calibrated in the 500-1000 Hz area, but does not correctly use the loop TF or the cavity pole.

As for the noise, remember that the whole point of changing the AS whitening gain was to turn on the whitening filter AFTER locking. With the WF OFF, there's no way that you can surpass the ADC noise limit.

Quote:

I calibrated the AS error signal into the displacement of the YARM cavity in the same way as I did before (elog).

No, I did not apply open loop TF to it (actually I could not measure the open loop TF because of the earthquake last night). So I should not have said it was the displacement.

Also I changed the AS gain with whitening filter on and xarm locked. Still it does not make any change.

  7832   Fri Dec 14 09:31:59 2012 ranaUpdateLSCHigh frequency noise in AS signal

This is NOT calibrated. Its sort of calibrated in the 500-1000 Hz area, but does not correctly use the loop TF or the cavity pole.

As for the noise, remember that the whole point of changing the AS whitening gain was to turn on the whitening filter AFTER locking. With the WF OFF, there's no way that you can surpass the ADC noise limit.

Quote:

I calibrated the AS error signal into the displacement of the YARM cavity in the same way as I did before (elog).

  7831   Fri Dec 14 09:05:04 2012 steveUpdatePEMTT after eq 6.1

 

 SRM, ITMY and MC3 moved some what, but how are the TT? It may worth looking at them before we vent.

  7830   Fri Dec 14 03:49:02 2012 AyakaUpdateGeneralEarthquake

There was an earthquake around 2:30 am. Now all the mirrors except SRM are damped.

earthquake.pngseis.pdf

  7829   Fri Dec 14 03:32:51 2012 AyakaUpdateLSCHigh frequency noise in AS signal

I calibrated the AS error signal into the displacement of the YARM cavity in the same way as I did before (elog).

The open loop transfer function is:

XARM_OL.pdf


The transfer function from ITMX excitation to AS error signal is:

AS_ITMXexc.pdf

Then I have got the calibration value : 5.08e+11 [counts/m]

The calibrated spectrum in unit of m/rtHz is

ASspe_noise.pdf

REF0: arm displacement
REF1: dark noise + demodulation circuit noise + WT filter noise + ADC noise (PSL shutter on)
REF2: demodulation circuit noise + WT filter noise + ADC noise (PD input of the circuit (at 1Y2) is connected to the 50 Ohm terminator)
(The circuit and WT filter seem to be connected at back side of the rack. Actually there is a connector labelled 'I MON' but it is not related to C1:LSC-ASS55_I_ERR)

Also we changed the AS gain so that ADC noise does not affect:

ASgain.png

However, this did not make big change in sensitivity. I guess this means that circuit noise limits the sensitivity at higher frequencies than 400 Hz.
I tried to adjust the AS gain carefully but I could not do that because of the earthquake. Further investigation is needed.

 

  7828   Fri Dec 14 03:15:49 2012 JenneUpdateSUSTT angle of incidence

But have you looked yet at what angle we need? The first input string mirror has a quite small incidence angle. The other input steering mirror maybe borderline, based on your estimates. Also, have we considered just having new brackets made and cleaned? Is there a reason we would prefer to modify the ones we have?

  7827   Fri Dec 14 02:52:32 2012 KojiUpdateSUSTT angle of incidence

At least, we don't want to use Al-coated mirrors. We should use multilayer dielectric mirrors.

  7826   Fri Dec 14 01:42:53 2012 DenUpdateSUSTT angle of incidence

I've estimated max possible angle of incidence on TT if we allow 20mm tolerance for the beam size and 5 mm tolerance for spot location on the mirror. It turns out to be

alpha = 43 degrees

So we need to cut the central part of the bracket. Then the max possible angle of incidence will be

alpha = 63 degrees

 

DSC_4791.jpeg

We can start the vent on Monday and use TT with an old bracket for yaw damping and later during the week we can install the brackets after they will be baked.

  7825   Thu Dec 13 21:21:34 2012 DenUpdatePEMseismometers

Quote:

 

 Looking at the PEM BLRMS, I noticed that the GUR1Z channel had a much reduced microseism compared to the GUR1X. Looking at the BLRMS screens everything seems ON, although its a mess (too many filters in the banks, etc. - clean this up, PEM people).

 Looking at the Z channel in DTT, I see that the Z spectra looks double high pass filtered below ~1 Hz.     Needs some attention in the daytime.

From Den and Ayaka's elog entry from Nov 29, its clear that this problem is there at that time. It seems that the seismometer was not even hooked up before then. Perhaps Tara returned the seismometer around Thanksgiving and then someone here hooked it up but neglected to log this work? If so, please make an elog now describing the installation of this sensor at the 40m and log any future work which takes place at the 40m lab even if you think it is inconsequential.

 Yesterday I wanted to recenter Guralps. I turned them off, understood that would be able to center them because we do not have power cable to Guralp box from Tara yet and turned them back on.

I've switched Guralp cables and spectrums are fine now.

  7824   Thu Dec 13 18:06:59 2012 AyakaUpdateWienerFilteringNew microphone mounts

 Yesterday, I made new mounts for microphones.

IMG_0129.JPGIMG_0130.JPG

I glued a microphone on a pedestal. The cables are attached loosely so that its tension does not make any noise.
At the bottom of the mount, I attached the surgical tube forming a ring by double-side tape so that it damps the seismic vibration.

I made 6 mounts and these are all on the AS table now.
I took some data of XARM signal controlled by AS.

My plan is to find/set an upper limit on acoustic coupling noise in AS signal.
The acoustic noise can be estimated by the Wiener filter, but it is not accurate because it may see residual correlation between AS and microphone signals that should be 0 when the data is long enough.
I will find/set an upper limit by the analysis based on Neuman-Pearson criterion, that is analog of a stochastic GW background search.
If I can find the acoustic coupling noise should be below the shot noise, I am happy. If not, some improvements may be needed someday.

  7823   Thu Dec 13 17:24:53 2012 AyakaUpdatePSLPMC calibration for MC_F calibration

 I calibrated MC_F signal into Hz/rtHz unit using the transfer function from MC_F to PMC feedback signal.

Here is the diagram:
  MC-PMCservo.jpg

n_mcf is MC_F signal we can get at dtt. I measured n_pmc/n'_mcf using SR.

TF_mcf-pmc.png

Other information I used:

G_out = 2.49/123.49 (see the document D980352-E01-C)

Fout has 1 pole at 10 Hz (see the document D980352-E01-C)

A_pzt = 371e+6/63 [Hz/V] (see elog)

F_wt has 1 pole at 100 Hz and 1 zero at 10 Hz.

Then, calibration transfer function of H is fitted as 1e+9/f [Hz/V]:

H.png

Then, the calibrated spectrum of MC_F is below:

 MCF_calib.pdf

This calibration have about 20 % error.
Compared to the spectrum in Jenne's paper (elog), above 20 Hz it seems to be laser frequency noise. But now we have extra unknown noise below 10 Hz.

Note: calibration value of laser's PZT is ~ 1MHz/V. This is reasonable compared to the data sheet of the laser. (This is calculated by combining result of H and transfer function of the circuit box1 and FSS.)

 laser.png

  7822   Thu Dec 13 04:42:32 2012 DenUpdateSUSITMX local damping

Tonight we've noticed that ITMX local damping was kicking the optics. This happened because LR shadow sensor was not working. In ~30 minutes it started to work again. Evan and I were working on installation, moving and focusing cameras and locking prcl and mich. We've installed a camera on BSC and plugged it in to PSL_SPARE input.

I'm not sure that this can be correlated to ITMX LR shadow sensor behaviour.

 

lrsen.png

  7821   Thu Dec 13 04:29:34 2012 DenUpdateSUSTT angle of incidence

I think the angle of incidence on TT inside BSC will be too large because of eddy current damping brackets. I've measured max possible angle of incidence

  Max angle of incidence, degrees
No bracket 72
Original bracket 45
New bracket (with no screws for tiny yaw magnets) 52

This means that we do not have too much range and there is a probability that 45 degree incident beam will start clipping. I think we should just cut off the central part of the bracket. We do not need it anyway, our eddy current damping due to corner magnets is good enough.

I've left the brackets near the laptop in the clean room.

  7820   Thu Dec 13 03:20:48 2012 DenUpdateLSCbeam inside DRMI is clipping on PR3 Tip-Tilt

Quote:

 

 I've made snapshots of PR2, PRM, ITMY and ITMX mirrors. Power buildup recycling gain (POWER BS / POWER PRM) was equal to 3-4.

           

 We've looked at PR2 face camera when PRM, BS and one of the ITMs were aligned. We saw an extra beam at PR2 when ITMX was aligned (right plot). This spot stays on the PR2 when prcl is locked.

PR2_ITMX.png   PR2_ITMY.png

Then we looked at PR3 transmission mirror and saw that the main beam is not on the edge of the mirror. Secondary beam is clipping on the mirror mount of PR3 that we see on BS_PRM camera.

PR3_LOCK.png

Measured beam spot positions:

Optics Pitch, mm Yaw, mm
ITMX 5.6 1.5
ETMX -1.5 1.5
ITMY 4.8 -1.5
ETMY -1.4 5.6
PRM 2.7 4.1

"+" for pitch means that the beam is too high, "-" too low

"+" for yaw means that the beam is left if you look from the back, "-" is right

Beam spots were measured using x, y arm and prcl locking to the carrier.

ELOG V3.1.3-